# Gravitational waves from cosmic strings in LISA: reconstruction pipeline and physics interpretation

### v1

@@ 2.4 Effects of non-standard cosmologies on the SGWB
The full SGWB spectrum can be obtained by summing over oscillation modes. We find that for $f\lesssim f_{\rm rd}$ the full SGWB spectrum is completely unaffected by this modification and may simply be computed as described in section 2.2 and in appendix A. For $f>f_{\rm rd}$, however, we find
@@
Figure 2: Examples of the full SGWB spectrum generated by cosmic strings in modified pre-BBN scenarios. The dashed lines represent the spectra generated by cosmic string evolving according to the SM of Cosmology and assuming the particle content of the SM of Particle Physics. Here we took $G\mu=10^{-10}$, $\alpha=10^{-1}$, $q=4/3$ and $\mathcal{F}=0.1$, and the shaded area corresponds to the LISA power-law sensitivity window. The left panel displays the results for various benchmark scenarios in which the universe underwent a period with a modified equation of state $w$ before the recent radiation dominated era, starting at $T_{\rm rd}=1\,\rm GeV$ (represented by the colored solid lines). The right panel shows the results with various benchmark numbers of new relativistic DoF, $\Delta g_{*}$, that decouple around $T_{\Delta}=10\,\rm MeV$ (the colored solid lines).
II. String network diluted by inflation. Modifications to the large-scale evolution of the cosmic string network may also induce modifications to the slope of the SGWB of the form discussed above. A particularly well-motivated example of such a scenario is that of cosmic strings created during the inflationary era Guedes:2018afo ; Gouttenoire:2019kij ; Cui:2019kkd . In this case, cosmic strings are diluted by the accelerated expansion of the universe and thus their correlation length is much larger than the horizon at the end of inflation. The network is then frozen in a non-relativistic conformal stretching regime after inflation and loop production is suppressed as a result. Significant loop production and GW emission only start once the cosmic string network re-enters the horizon and attains relativistic velocities and, as a result, the SGWB is suppressed at large frequencies. In this case, the fundamental mode of the spectrum exhibits a characteristic $f^{-1}$ cut-off in this frequency range Guedes:2018afo , corresponding to a slope of $q-1$ in the full SGWB Cui:2019kkd . The frequency where this cut-off starts is determined by the value of the scale factor at the time when the strings re-enter the Hubble volume, $a_{e}$, and given by Guedes:2018afo
@@
where $\epsilon_{r}=\alpha/(\Gamma G\mu)$ and we have assumed that this happens before the end of the radiation era, or equivalently for $f_{\rm rd}>\left(9.4\times 10^{-18}/G\mu\right)\,{\rm Hz}$ (see ref. Guedes:2018afo for the complete expression). The modification to the spectrum caused by this deviation from the standard evolution of cosmic string networks is then also of the form of eq. (12), with $d=1$. In this case, the predicted spectrum is then also identical to that predicted for the scenarios in which there is a modified equation of state with $w\leq(3-q)/(q+1)/3$, which means that these two scenarios may be indistinguishable  [^13] Except for the fact that, in this case, the spectrum may be modified up to lower frequencies since cosmic strings may re-enter the horizon at late cosmological times..
III. New particle species. Another type of well-motivated modification to the standard scenario is the inclusion of additional particle species that are relativistic in the early universe and contribute to $g_{*}$ and $g_{*S}$. Such new states are ubiquitously predicted in the extensions to the SM. In some cases, the increase in $g_{*}$ can be dramatic: for instance, the minimal supersymmetric extension of the SM predicts roughly a hundred new DoF beyond the $g_{*}=106.75$ present in the SM, and in mirror dark sector scenario the number of new DOFs equals that of the SM Kolb:1985bf ; Berezhiani:1995am .
@@
The precise evolution of $g_{*},g_{*s}$ depends on the specific dynamics of the underlying Beyond the Standard Model (BSM) theory. To illustrate the generic effect of new relativistic DoFs on the string GW spectrum, we take a simple approach and model the change as a rapid decrease in $g_{*}$ as the temperature falls below the threshold $T_{\Delta}$  [^14] For new DoFs coupled to the SM plasma, $T_{\Delta}$ sets the temperature scale at which the new particles are Boltzmann suppressed, i.e. their mass scale. In the presence of dark sectors, the associated dark plasma may have its own temperature independent of that of the SM. In this case, the connection between $T_{\Delta}$ and the mass scale of the new DoFs is loose.:
@@
where the $f_{\Delta}$-$T_{\Delta}$ relation is the same as the $f_{\rm rd}$-$T_{\rm rd}$ as defined in eq. (11). Equation (15) provides a good analytic estimate for the decrease in the magnitude of $\Omega_{\rm gw}$ in the high frequency plateau. However, depending on $T_{\Delta}$, the LISA band may not be able to capture such a plateau region. Instead, the transitional region may lie within LISA reach. The spectral shape for this transitional region is more subtle, as it depends on $T_{\Delta}$ and the details of how these new DoFs enter/decouple from the radiation thermal bath. Moreover, the sudden change in the expansion rate caused by this decrease has a non-linear impact on the large-scale dynamics of cosmic string networks, which is not straightforward to model but should also affect the shape of the spectrum in this transitional region. Here, for simplicity, we assume that all additional DoFs decouple instantaneously around the same temperature $T_{\Delta}$ and that cosmic strings remain in a linear scaling regime during the process. The templates for Models I and II are then recalculated, using the same methods described in the previous sections, for this modified thermal history. The impact of these additional DoFs on the cosmic SGWB is illustrated in the right panel of figure 2.
@@ 3 Template-based reconstruction with the SGWBinner pipeline
In this work, as well as in the papers Caprini:2024hue ; LISACosmologyWorkingGroupInflation , we introduce the option for template-based analysis of a primordial SGWB signal within SGWBinner. This paper’s focus is on cosmic string SGWBs  [^15] The same development is presented in refs. Caprini:2024hue ; LISACosmologyWorkingGroupInflation for the implementation of the first-order phase transition and inflationary templates. . To elucidate this new option, we provide a brief overview of the foundational aspects of the code and LISA measurements.
@@ 3.1 LISA measurement channels
Within the above approximations, casting the phase measurements in terms of the channels A, E, and T gives important advantages as they virtually constitute three independent GW interferometers  [^16] See, e.g., Hartwig:2023pft for an analysis of how the non-orthogonality of TDI variables, induced by non-equilateral constellations and unequal noise levels, affects signal parameter reconstruction.. The A and E channels exhibit identical properties while the sensitivity of channel T to the signal is significantly lower and it can be treated as a quasi-null channel Prince:2002hp . Due to these properties, the SGWBinner code generates mock data by simulating these three channel’s measurements.
@@ 3.3.3 Galactic foreground
$h^{2}\Omega^{\textrm{Gal}}_{\rm gw}(f)=h^{2}\Omega_{\rm Gal}\frac{f^{3}}{2}\left(\frac{f}{\textrm{Hz}}\right)^{-\frac{7}{3}}\left[1+\tanh\left({\frac{f_{\textrm{knee}}-f}{f_{2}}}\right)\right]e^{-(f/f_{1})^{\upsilon}}\;,$ (26)
@@ 3.3.4 Cosmic string SGWB
Parameter name Parameter symbol Range
String tension (Model I) $\log_{10}(G\mu)$ $[-18.0,-7.0]$
String tension (Model II) $\log_{10}(G\mu)$ $[-18.0,-9.5]$
@@
Loop size $\log_{10}(\alpha)$ $[-3,0]$
New degrees of freedom (DoF) $\Delta g$ $[0,120]$
Temperature of new DoF decoupling $\log_{10}(T_{\Delta}/1\,\text{GeV})$ $[-2.3,1.7]$
Equation-of-state $w$ $[0,1]$
Temperature of radiation domination $\log_{10}(T_{rd}/1\,\text{GeV})$ $[-2.3,1.7]$
Table 1: The parameters used in various cosmic string signal models, along with their associated ranges. All parameters listed here are drawn from flat priors during reconstruction.
@@ 4.1 Results for Model I
For Model I, we use the semi-analytic template discussed in section 2.2. Its free parameters are $G\mu$, $\alpha$, and $q$, although other parameters may also be included [^18] For example, in section 4.4, we also include the “fuzziness” parameter, $\mathcal{F}$, which stands in for several effects, such as the amount of kinetic energy carried out by the loops produced by the network as well as the spread in size.. For the fuzziness parameter, we set $\mathcal{F}=0.1$.
@@
Figure 4 presents the exploration of parameter space for 1 $\sigma$ relative errors in the $\log_{10}G\mu-q$ space, with $\log_{10}\alpha$ fixed at 1.01.0-1.0- 1.0.191919For lower values of $\alpha$, the amplitude of the SGWB is lower and, as a result, the lowest tension that LISA may probe is increased (see, e.g., ref. Auclair:2019wcv ). For a given parameter, the relative error is derived from diagonal entry in the (inversed) reconstructed covariance matrix; this corresponds to the error on the given parameter after marginalizing over all the others. In the figure we display the cases with and without foregrounds, as the behavior related to the cosmic string model is more prominent in the former case, while the latter represents a more realistic situation. [^20] Possibly, by leveraging the yearly modulation of the galactic foreground, one could obtain an intermediate result. We observe a general trend where the error is smallest at intermediary values of the tension (between $G\mu\sim 10^{-16}$ and $10^{-13}$). The increase in error at low tension ($G\mu<10^{-17}$) is easily explained by the SGWB signal dropping much below the LISA power-law sensitivity curve. On the other hand, at large tensions, the increase in error is due to correlations in the parameters.
@@ 4.3 Probing the pre-BBN expansion rate with reconstruction
Unsurprisingly, the quality of the reconstruction depends strongly on the values of $\Delta g$ and $T_{\Delta}$. The result for the Model II signal with $\log_{10}(G\mu)=-10.0$, $\Delta g=10$, and $T_{\Delta}=10^{-1.5}\,\text{GeV}$ in the presence of foregrounds is shown in figure 10. In this case, the $\Delta g=0$ is significantly disfavored, and thus this cosmic string signal would allow us to also confirm the presence of new DoF from beyond the SM. We also test the reconstruction of this same new DoF setting in the case of Model I signal in figure 11. We observe that the reconstruction is more challenging, as the injected value of $\Delta g=10$ is compatible with $\Delta g=0$ at 2 $\sigma$. The degradation in the reconstruction is partly due to the two additional free parameters involved in Model I. This model also includes the evolution of the loop number density due to the modified expansion. Given the relatively slow relaxation of the network, the modification is then spread over a longer time, resulting in a smoother spectrum which also makes the reconstruction more challenging. While finding the imprint of new degrees of freedom is more difficult, in the spectra of Model I it is not impossible. In figure 12 we show that a successful reconstruction is possible with the number of new degrees of freedom increased to $\Delta g=50$.
Figure 10: The reconstruction of a Model II cosmic string SGWB, in the presence of foregrounds (not shown in the corner plots), at $\log_{10}(G\mu)=-10.0$ with modifications due to $10$ new DoF added at $10^{-1.5}\,\text{GeV}$. Repeated fits of the same injected signal consistently produce closed contours that contain the correct values, although the biases of the returned means can change.
Figure 11: The reconstruction of a Model I cosmic string SGWB, in the presence of foregrounds, at $\log_{10}(G\mu)=-10.0$ with modifications due to $10$ new DoF added at $10^{-1.5}\,\text{GeV}$.
Figure 12: The reconstruction of a Model I cosmic string SGWB, in the presence of foregrounds, at $\log_{10}(G\mu)=-10.0$ with modifications due to $50$ new DoF added at $10^{-1.5}\,\text{GeV}$.
@@
In figure 13, we show an example of the parameter estimation for Model I with two additional free parameters: the equation-of-state parameter $w$, which is set to 1 assuming a kination phase, and the temperature of the universe when the non-standard evolution ends and switches to the radiation-dominated universe, denoted as $T_{rd}$. The reconstruction is quite successful in this example, but, as before, the quality of the results will necessarily be better for lower $T_{rd}$ and will also depend strongly on the value of $w$. In particular, as shown in section 4.3, for $w<1/3$ the spectrum develops a negative slope for $f>f_{rd}$ and its amplitude quickly drops below that of the foregrounds and below the LISA sensitivity window. Also, as previously discussed, for all $w\leq 1/9$ the predicted spectrum is the same and very similar to that of strings created during inflation, which makes the reconstruction of $w$ more challenging.
Figure 13: The reconstruction of a Model I cosmic string SGWB, in the presence of foregrounds, at $\log_{10}G\mu=-10.0$ with non-standard equation of state $w=1$.

### v2

@@ 2.4.1 Modified equation of state
The full SGWB spectrum may again be obtained by summing over harmonic modes. We find that for $f\lesssim f_{\rm rd}$ the full SGWB spectrum is completely unaffected by this modification and may simply be computed as described in section 2.2 and in appendix A. For $f>f_{\rm rd}$, however, we find
@@
Figure 2: Examples of the full SGWB spectrum generated by cosmic strings in modified pre-BBN scenarios. The dashed lines represent the spectra generated by cosmic string evolving according to the SM of Cosmology and assuming the particle content of the SM of Particle Physics. Here we took $G\mu=10^{-10}$, $\alpha=10^{-1}$, $q=4/3$ and $\mathcal{F}=0.1$, and the shaded area corresponds to the LISA power-law sensitivity window. The left panel displays the results for various benchmark scenarios in which the universe underwent a period with a modified equation of state $w$ before the recent radiation dominated era (Model Ia), starting at $T_{\rm rd}=1\,\rm GeV$ (represented by the colored solid lines). The right panel shows the results with various benchmark numbers of new relativistic DoF, $\Delta g_{*}$, based on Model Ib, that decouple around $T_{\Delta}=10\,\rm MeV$ (the colored solid lines).
2.4.2 String network diluted by inflation
Modifications to the large-scale evolution of the cosmic string network may also induce modifications to the slope of the SGWB of the form discussed above. A particularly well-motivated example of such a scenario is that of cosmic strings created during the inflationary era Shafi:1984tt ; Vishniac:1986sk ; Yokoyama:1988zza ; Yokoyama:1989pa ; Jeannerot:2003qv (for a review of later works, see Auclair:2019wcv ; LISACosmologyWorkingGroup:2022jok ). In this case, cosmic strings are diluted by the accelerated expansion of the universe and thus their correlation length is much larger than the horizon at the end of inflation. The network is then frozen in a non-relativistic conformal stretching regime after inflation and loop production is suppressed as a result. Significant loop production and GW emission only start once the cosmic string network re-enters the horizon and attains relativistic velocities and, as a result, the SGWB is suppressed at large frequencies Guedes:2018afo ; Gouttenoire:2019kij ; Cui:2019kkd . In this case, the fundamental mode of the spectrum exhibits a characteristic $f^{-1}$ cut-off in this frequency range Guedes:2018afo , corresponding to a slope of $q-1$ in the full SGWB Cui:2019kkd . The frequency where this cut-off starts is determined by the value of the scale factor at the time when the strings re-enter the Hubble volume, $a_{e}$, and given by Guedes:2018afo
@@ 2.4.2 String network diluted by inflation
where $\epsilon_{r}=\alpha/(\Gamma G\mu)$ and we have assumed that this happens before the end of the radiation era, or equivalently for $f_{\rm rd}>\left(9.4\times 10^{-18}/G\mu\right)\,{\rm Hz}$ (see ref. Guedes:2018afo for the complete expression). The modification to the spectrum caused by this deviation from the standard evolution of cosmic string networks is then also of the form of eq. (12), with $d=1$. In this case, the predicted spectrum is then also identical to that predicted for the scenarios in which there is a modified equation of state with $w\leq(3-q)/(q+1)/3$, which means that these two scenarios may be indistinguishable. [^12] Except for the fact that, in this case, the spectrum may be modified up to lower frequencies since cosmic strings may re-enter the horizon at late cosmological times. Template Ia may then also be used to probe this scenario by setting $w\leq(3-q)/(q+1)/3$ and, in this case, the only additional free parameter would be $a_{e}$.
2.4.3 New particle species
Another type of well-motivated modification to the standard scenario is the inclusion of additional particle species that are relativistic in the early universe and contribute to $g_{*}$ and $g_{*S}$. Such new states are ubiquitously predicted in the extensions to the SM. In some cases, the increase in $g_{*}$ can be dramatic: for instance, the minimal supersymmetric extension of the SM predicts roughly a hundred new DoF beyond the $g_{*}=106.75$ present in the SM, and in mirror dark sector scenario the number of new DOFs equals that of the SM Kolb:1985bf ; Berezhiani:1995am .
@@ 2.4.3 New particle species
The precise evolution of $g_{*},g_{*s}$ depends on the specific dynamics of the underlying Beyond the Standard Model (BSM) theory. To illustrate the generic effect of new relativistic DoFs on the string GW spectrum, we take a simple approach and model the change as a rapid decrease in $g_{*}$ as the temperature falls below the threshold $T_{\Delta}$: [^13] For new DoFs coupled to the SM plasma, $T_{\Delta}$ sets the temperature scale at which the new particles are Boltzmann suppressed, i.e. their mass scale. In the presence of dark sectors, the associated dark plasma may have its own temperature independent of that of the SM. In this case, the connection between $T_{\Delta}$ and the mass scale of the new DoFs is loose.
@@
where the $f_{\Delta}$-$T_{\Delta}$ relation is the same as the $f_{\rm rd}$-$T_{\rm rd}$ as defined in eq. (11). Equation (15) provides a good analytic estimate for the decrease in the magnitude of $\Omega_{\rm gw}$ in the high frequency plateau. However, depending on $T_{\Delta}$, the LISA band may not be able to capture such a plateau region. Instead, the transitional region may lie within LISA reach. The spectral shape for this transitional region is more subtle, as it depends on $T_{\Delta}$ and the details of how these new DoFs enter/decouple from the radiation thermal bath. Moreover, the sudden change in the expansion rate caused by this decrease has a non-linear impact on the large-scale dynamics of cosmic string networks, which is not straightforward to model but should also affect the shape of the spectrum in this transitional region. Here, for simplicity, we assume that all additional DoFs decouple instantaneously around the same temperature $T_{\Delta}$ and that cosmic strings remain in a linear scaling regime during the process. The templates for Models I and II are then recalculated, using the same methods described in the previous sections, for this modified thermal history to develop Models Ib and IIb, which now include two additional free parameters ($T_{\Delta}$ and $\Delta g_{*}$). Model I as always involves the analytical description of the evolution of the network throughout the transition. At the same time in Model II we assume an instantaneous relaxation of the network and the effect of the modification comes mostly through modified redshift of the signal. The impact of these additional DoFs on the cosmic SGWB from Model I is illustrated in the right panel of figure 2.
@@ 3 Template-based reconstruction with the SGWBinner pipeline
In this work, as well as in the papers Caprini:2024hue ; LISACosmologyWorkingGroupInflation , we introduce the option for template-based analysis of a primordial SGWB signal within SGWBinner. This paper’s focus is on cosmic string SGWBs. [^14] The same development is presented in refs. Caprini:2024hue ; LISACosmologyWorkingGroupInflation for the implementation of the first-order phase transition and inflationary templates. To elucidate this new option, we provide a brief overview of the foundational aspects of the code and LISA measurements.
@@ 3.1 LISA measurement channels
Within the above approximations, casting the phase measurements in terms of the channels A, E, and T gives important advantages as they virtually constitute three independent GW interferometers. [^15] See, e.g., Hartwig:2023pft for an analysis of how the non-orthogonality of TDI variables, induced by non-equilateral constellations and unequal noise levels, affects signal parameter reconstruction. The A and E channels exhibit identical properties while the sensitivity of channel T to the signal is significantly lower and it can be treated as a quasi-null channel Prince:2002hp . Due to these properties, the SGWBinner code generates mock data by simulating these three channel’s measurements.
@@ 3.3.3 Galactic foreground
$h^{2}\Omega^{\textrm{Gal}}_{\rm gw}(f)=\frac{1}{2}h^{2}\Omega_{\rm Gal}\left(\frac{f}{\textrm{Hz}}\right)^{\frac{2}{3}}\left[1+\tanh\left({\frac{f_{\textrm{knee}}-f}{f_{2}}}\right)\right]e^{-(f/f_{1})^{\upsilon}}\;,$ (26)
@@ 3.3.4 Cosmic string SGWB
Scenario Parameter name Symbol Range
Model I String tension $\log_{10}(G\mu)$ $[-18.0,-7.0]$
Loop size $\log_{10}(\alpha)$ $[-3,0]$
@@
Fuzziness $\log_{10}(\mathcal{F})$ $[-5,1]$
[…]
Rad. domination temp. $\log_{10}(T_{rd}/\text{GeV})$ $[-2.3,1.7]$
[…]
Decoupling temp. $\log_{10}(T_{\Delta}/\text{GeV})$ $[-2.3,1.7]$
Model II String tension $\log_{10}(G\mu)$ $[-18.0,-9.5]$
[…]
Decoupling temp. $\log_{10}(T_{\Delta}/\text{GeV})$ $[-2.3,1.7]$
Table 1: The parameters used in reconstructing signals in sec. 4, along with their associated ranges. All parameters listed here are drawn from flat priors during reconstruction. Note that Model I’s fuzziness parameter is set to $\mathcal{F}=0.1$ in all scenarios except for the reconstructions of sec. 4.5.
@@ 4.1 Results for Model I
For Model I, we use the semi-analytic template discussed in section 2.2. Its free parameters are $G\mu$, $\alpha$, and $q$, although other parameters may also be included. [^17] For example, in section 4.4, we also include the “fuzziness” parameter, $\mathcal{F}$, which stands in for several effects, such as the amount of kinetic energy carried out by the loops produced by the network as well as the spread in size. Here, we will follow the prescription in Auclair:2019wcv to study Nambu-Goto strings and take $\mathcal{F}=0.1$ as the standard normalization, as this allows us to match the amplitude of the loop number density inferred from the Nambu–Goto simulations of ref. Blanco-Pillado:2013qja .
@@
Figure 4 presents the exploration of parameter space for 1 $\sigma$ relative errors in the $\log_{10}G\mu-q$ space, with $\log_{10}\alpha$ fixed at 1.01.0-1.0- 1.0.181818For lower values of $\alpha$, the amplitude of the SGWB is lower and, as a result, the lowest tension that LISA may probe is increased (see, e.g., ref. Auclair:2019wcv ). For a given parameter, the relative error is derived from diagonal entry in the (inversed) reconstructed covariance matrix; this corresponds to the error on the given parameter after marginalizing over all the others. In the figure we display the cases with and without foregrounds, as the behavior related to the cosmic string model is more prominent in the former case, while the latter represents a more realistic situation. [^19] Possibly, by leveraging the yearly modulation of the galactic foreground, one could obtain an intermediate result. We observe a general trend where the error is smallest at intermediary values of the tension (between $G\mu\sim 10^{-16}$ and $10^{-13}$). The increase in error at low tension ($G\mu<10^{-17}$) is easily explained by the SGWB signal dropping much below the LISA power-law sensitivity curve. On the other hand, at large tensions, the increase in error is due to correlations in the parameters.
@@ 4.3 Probing the pre-BBN expansion rate with reconstruction
Unsurprisingly, the quality of the reconstruction depends strongly on the values of $\Delta g$ and $T_{\Delta}$. The result for the Model IIb signal with $\log_{10}(G\mu)=-10.0$, $\Delta g=10$, and $T_{\Delta}=10^{-1.5}\,\text{GeV}$ in the presence of foregrounds is shown in figure 10. In this case, the $\Delta g=0$ is significantly disfavored, and thus this cosmic string signal would allow us to also confirm the presence of new DoF from beyond the SM. We also test the reconstruction of this same new DoF setting in the case of Model Ib signal in figure 11 (with $\mathcal{F}=0.1$ as well). We observe that the reconstruction is more challenging, as the injected value of $\Delta g=10$ is compatible with $\Delta g=0$ at 2 $\sigma$. The degradation in the reconstruction is partly due to the two additional free parameters involved in Model I. This model also includes the evolution of the loop number density due to the modified expansion. Given the relatively slow relaxation of the network, the modification is then spread over a longer time, resulting in a smoother spectrum which also makes the reconstruction more challenging. While finding the imprint of new degrees of freedom is more difficult, in the spectra of Model I it is not impossible. In figure 12 we show that a successful reconstruction is possible with the number of new degrees of freedom increased to $\Delta g=50$.
Figure 10: The reconstruction of a Model IIb cosmic string SGWB, in the presence of foregrounds (not shown in the corner plots), at $\log_{10}(G\mu)=-10.0$ with modifications due to $10$ new DoF added at $10^{-1.5}\,\text{GeV}$. Repeated fits of the same injected signal consistently produce closed contours that contain the correct values, although the biases of the returned means can change.
Figure 11: The reconstruction of a Model Ib cosmic string SGWB, in the presence of foregrounds, at $\log_{10}(G\mu)=-10.0$ with modifications due to $10$ new DoF added at $10^{-1.5}\,\text{GeV}$.
Figure 12: The reconstruction of a Model Ib cosmic string SGWB, in the presence of foregrounds, at $\log_{10}(G\mu)=-10.0$ with modifications due to $50$ new DoF added at $10^{-1.5}\,\text{GeV}$.
@@
In figure 13, we show an example of the parameter estimation for Model Ia, with its two additional free parameters: the equation-of-state parameter $w$, which is set to 1 assuming a kination phase, and the temperature of the universe when the non-standard evolution ends and switches to the radiation-dominated universe, denoted as $T_{rd}$ (we also took $\mathcal{F}=0.1$). The reconstruction is quite successful in this example, but, as before, the quality of the results will necessarily be better for lower $T_{rd}$ and will also depend strongly on the value of $w$. In particular, as shown in section 4.3, for $w<1/3$ the spectrum develops a negative slope for $f>f_{rd}$ and its amplitude quickly drops below that of the foregrounds and below the LISA sensitivity window. Also, as previously discussed, for all $w\leq 1/9$ the predicted spectrum is the same and very similar to that of strings created during inflation, which makes the reconstruction of $w$ more challenging.
Figure 13: The reconstruction of a Model Ia cosmic string SGWB, in the presence of foregrounds, at $\log_{10}G\mu=-10.0$ with non-standard equation of state $w=1$.